OpenStudy (anonymous):

Need help with this homogeneous equation dy/dx= (x^2-y^2)/3xy.

OpenStudy (anonymous):

put y=vx \[\frac{ dy }{ dx }=v+x \frac{ dv }{ dx }\]

OpenStudy (anonymous):

how did you get that? what did you substitute exactly?

OpenStudy (freckles):

http://tutorial.math.lamar.edu/Classes/DE/Substitutions.aspx It is a common substitution.

OpenStudy (freckles):

have you tried that sub yet?

OpenStudy (freckles):

instead of in terms of y and x you will have a differential equation in terms of v and x

OpenStudy (anonymous):

I am trying something else out and I think I got it. The notation threw me off.

OpenStudy (freckles):

what did you try?

OpenStudy (anonymous):

never mind the integral was really ugly I tried multiplying by (y/x^3) to get in terms of (y/x)

OpenStudy (freckles):

ok well v=y/x can be written as vx=y (by multiply x on both sides) wherever you see y put vx but wherever you see dy/dx put v'x+v

OpenStudy (freckles):

to get from y=vx to y'=v'x+v you just need product rule

OpenStudy (freckles):

ok so we have dy/dx= (x^2-y^2)/3xy. so replace dy/dx with y'=v'x+v and replacing y with vx we have \[v'x+v=\frac{x^2-(vx)^2}{3x(vx)}\]

OpenStudy (freckles):

try to simplify that one fraction first

OpenStudy (anonymous):

so I should get v'x+v=(x^2-(vx)^2)/3x(vx)

OpenStudy (anonymous):

sweet got that

OpenStudy (freckles):

you should hopefully realize you have a separable differential equation

OpenStudy (freckles):

after of course simplifying that one fraction :)

OpenStudy (anonymous):

When I simplified I got 1/3z - z/3 is that right? Want to make sure my algebra is correct

OpenStudy (anonymous):

whoops 1/3v-v/3

OpenStudy (freckles):

\[\frac{x^2-(vx)^2}{3x(vx)}=\frac{x^2-v^2x^2}{3x^2v}=\frac{x^2(1-v^2)}{x^2(3v)} \\ =\frac{\cancel{x^2}(1-v^2)}{\cancel{x^2}(3v)}=\frac{1-v^2}{3v}\] yep that looks like what you got you just wrote it as two separate fractions

OpenStudy (freckles):

\[v'x+v=\frac{1-v^2}{3v}\] \[\frac{dv}{dx}x+v=\frac{1-v^2}{3v} \\ \frac{dv}{dx} x =\frac{1-v^2}{3v}-v \] \[\frac{dv}{dx}x=\frac{1-4v^2}{3v}\] do you that we have a separable differential equation?

OpenStudy (freckles):

i will leave the separating to you but i will still be here if you have problems of course

OpenStudy (anonymous):

awesome! Thank you so much! I should be able to do the separating. Its the simplifying I get stuck at :)

OpenStudy (anonymous):

it is a homogeneous differential equation. To solve it is best substitution.